OpenStudy (bj11684):

A sports club has 21 basketball players. If 12% of the total number of members in the club are basketball players, what is the total number of members in the club? 120 135 170 175

OpenStudy (bj11684):

ok

OpenStudy (bj11684):

it did not work

OpenStudy (bj11684):

the answer choices it was wrong

OpenStudy (anonymous):

21/12%

OpenStudy (anonymous):

\[21\div \frac{ 12 }{ 100 }\]

OpenStudy (anonymous):

what is your answer

OpenStudy (anonymous):

are u with me

OpenStudy (anonymous):

its 0.0175

OpenStudy (anonymous):

so then its 175

OpenStudy (anonymous):

no that is not correct

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

when solving a solution like this you have to change the sign to multiplication which makes you swap the fraction by the right of the solution

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

now follow the hint i just throw up

OpenStudy (anonymous):

so then its 21\[21\times\frac{ 12 }{ 100}\] so

OpenStudy (anonymous):

or i swich it

OpenStudy (anonymous):

oh i swap the right

OpenStudy (anonymous):

then swap 12/100 with each other

OpenStudy (anonymous):

oh so 100/12

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[12\times \frac{ 100 }{ 12}\]

OpenStudy (anonymous):

or is it 21times 100/12

OpenStudy (anonymous):

@iseeunicorns_

OpenStudy (anonymous):

that is correct @Dahaba2.0

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

Okay have u seen kelly's mistake @iseeunicorns_

OpenStudy (anonymous):

wdym

OpenStudy (anonymous):

okay now what is your answer

OpenStudy (anonymous):

oh wow i said that a long time a go well at least ik how to show mt work and kelly neede to put the negtive sign in front of the 8

OpenStudy (anonymous):

i knew u gave a guess but i just wanted u to understand it

OpenStudy (anonymous):

oh ok lol nice